# Computing the Topological Degree of Maps Between 2-Spheres

### v3

@@ 2. Proofs
The function $T$ has no zeros on the unit circle. Suppose, for contradiction, that $T(e^{i\alpha_{0}})=0$ for some $\alpha_{0}\in[0,2\pi]$. Since $T$ is not identically zero, there exists a point $a$ such that $T(a)\neq 0$. By the homogeneity of $T$, the function $T(z)$ has no limit as $|z|\to\infty$: along the ray from the origin through $e^{i\alpha_{0}}$ the function vanishes identically, whereas along the ray through $a$ we have
$|T(ka)|=k^{d}|T(a)|\to\infty\quad\text{as }k\to\infty\text{ with }k>0.$
Therefore, the minimum of the modulus of $T$ on the unit circle is greater than 0. It then follows from the homogeneity of $T$ that there exists a sufficiently large number $M>0$ such that for all $|z|>M$ the numerator of $R$ has no zeros and
|T(z)|>|f(z)T(z)|.|T(z)|>|f(z)-T(z)|.
@@
We factor $\tilde{T}(z)$ as
(2) $\tilde{T}(z)=a_{d}\prod_{t=1}^{d}(z-z_{t}),$
where $z_{1},\ldots,z_{d}$ are the roots of $\tilde{T}(z)$ and $a_{d}$ is its leading coefficient.
@@
Thus,
$\text{ind}_{0}(\tilde{f})=\text{ind}_{0}(T(Me^{i\alpha}))=\text{ind}_{0}(M^{d}T(e^{i\alpha}))=\text{ind}_{0}(T(e^{i\alpha}))=\text{ind}_{0}(e^{-id\alpha}\tilde{T}(e^{i\alpha}))=\text{ind}_{0}(\tilde{T}(e^{i\alpha}))-d.$
The factors of $\tilde{T}$ in Equation (2) can be continuously deformed in $\mathbb{C}\setminus\{0\}$ without changing the winding number. Specifically, if a root $z_{t}$ has modulus greater than 1, the loop
$0,2\pi]\longrightarrow\mathbb{C}\setminus\{0\},\quad\alpha\longmapsto e^{i\alpha}-z_{t}$
is homotopic to a constant loop. If the modulus of a root $z_{t}$ is less than 1, it is homotopic to the loop
$0,2\pi]\longrightarrow\mathbb{C}\setminus\{0\},\quad\alpha\longmapsto e^{i\alpha}.$
Then, the index of the following curve:
$S^{1}\rightarrow S^{1}:\alpha\mapsto\tilde{T}(e^{i\alpha})$
is equal to the number of roots of the polynomial $\tilde{T}$ in the unit disk, counted with multiplicity. $\Box$

### v4

@@ 2. Proofs
The function $T$ has no zeros on the unit circle. Suppose, for contradiction, that $T(e^{i\alpha_{0}})=0$ for some $\alpha_{0}\in[0,2\pi]$. Since $T$ is homogeneous, we must have $T(z)=0$ for all $z$ on the ray through $e^{i\alpha_{0}}$. This contradicts the existence of an infinite limit of $T$ at infinity. Therefore, the minimum of the modulus of $T$ on the unit circle is greater than 0. It then follows from the homogeneity of $T$ that there exists a sufficiently large number $M>0$ such that for all $|z|>M$, the function $f(z)$ has no zeros, and
$\left|T(z)\right|>\left|f(z)-T(z)\right|.$
@@
Due to the homogeneity of $T$ and the fact that the winding number is invariant under scaling by a non-zero constant, we have:
$\text{ind}_{0}(T(Me^{i\alpha}))=\text{ind}_{0}(M^{d}T(e^{i\alpha}))=\text{ind}_{0}(T(e^{i\alpha}))$
@@
Applying the property that the index of a product of curves is the sum of their indices, we obtain:
$\text{ind}_{0}(\tilde{f})=\text{ind}_{0}(e^{-id\alpha}\tilde{T}(e^{i\alpha}))=\text{ind}_{0}(e^{-id\alpha})+\text{ind}_{0}(\tilde{T}(e^{i\alpha}))=-d+\text{ind}_{0}(\tilde{T}(e^{i\alpha}))$
Then, by the argument principle, the index of the curve $\tilde{f}$ is equal to the number of roots of the polynomial $\tilde{T}$ in the open unit disk, counted with multiplicity, minus $d$.
$\Box$